OpenStudy (kitkat1):

For f(x) = 0.01(2)x, find the average rate of change from x = 12 to x = 15. (2 points) 3 40.96 95.57 286.72

OpenStudy (kitkat1):

Plx help

OpenStudy (superdavesuper):

is it \[f(x) = 0.01(2^{x})\] or \[f(x)=0.01(2)x\] the two are quite different

OpenStudy (kitkat1):

Its like the first one

OpenStudy (superdavesuper):

alright that makes more sense :) u can find the ans by using the formula: average rate of change of f(x) between x=x1 and x=x2 : \[\frac{ f(x1) - f(x2) }{ x1 - x2 }\] here x1=12 and x2=15

OpenStudy (kitkat1):

How do I find the f(x1) and f(x2)?

OpenStudy (superdavesuper):

use the given formula for f(x)

OpenStudy (kitkat1):

95.57?

OpenStudy (superdavesuper):

what is equal to 95.7?

OpenStudy (kitkat1):

Ye